In a bumper test, three types of autos were deliberately crashed into a barrier at 5 mph, and the resulting damage (in dollars) was estimated. Five test vehicles of each type were crashed, with the results shown below.
Crash Damage ($)
Goliath Varmint Weasel
1,600 1,290 1,090
760 1,400 2,100
880 1,390 1,830
1,950 1,850 1,250
1,220 950 1,920
Research question: Are the mean crash damages the same for these three vehicles?
show work

Answer by checkley77(12569)

Goliath Varmint Weasel
1,600 1,290 1,090
760 1,400 2,100
880 1,390 1,830
1,950 1,850 1,250
1,220 950 1,920
---------- ---------------- ----------------
6,410/5=1,282 6,880/5=1,375 8,190/5=1,638
AS YOU CAN SEE THE MEANS (AVERAGES) ARE NOT THE SAME.